OpenStudy (egbeach):

Determine whether the vectors u and v are parallel, orthogonal, or neither. u = <6, -2>, v = <8, 24>

OpenStudy (usukidoll):

the vectors are orthogonal if the result is 0 after using the dot product. \[u \cdot v =<u_1v_1,u_2v_2,...u_nv_n> \]

OpenStudy (anonymous):

If vectors are parallel, then there exists a scalar \(c\) such that\[ \mathbf u = c \mathbf v \]

OpenStudy (usukidoll):

\[u \cdot v = <u_1v_1+u_2v_2+...u_nv_n>\] for dot product but if it's neither than the dot product can't be 0 and a scalar c doesn't exist.

OpenStudy (egbeach):

do i just plug in the numbers to solve?

OpenStudy (usukidoll):

you could.. since we have \[u=<u_1,u_2>, v=<v_1,v_2>\] and we can use dot product \[u \cdot v = <u_1v_1+u_2v_2+...u_nv_n> \]

OpenStudy (usukidoll):

u= <6,-2> and v = <8,24> \[u_1=6,u_2=-2,v_1=8,v_2=24 \] now we plug those values into the dot product formula \[u \cdot v = u_1v_1+u_2v_2\]

OpenStudy (usukidoll):

\[u \cdot v = (6)(8)+(-2)(24) \]

OpenStudy (usukidoll):

so what is 6 x 8 and what is -2 x 24 ?

OpenStudy (egbeach):

6x8=48 -2x24= -48

OpenStudy (usukidoll):

mhm \[u \cdot v = 48-48 \] so now what's 48-48?

OpenStudy (egbeach):

0. so is it neither?

OpenStudy (usukidoll):

no.. since our dot product is 0, our vectors are _______________

OpenStudy (egbeach):

orthogonal since it equal 0

OpenStudy (usukidoll):

yes :)